# TACO-RL: Task Aware Prompt Compression Optimization with Reinforcement Learning

### v2

@@ Appendix A Details of Experiments
For this dataset only, we used target summaries generated using GPT-4 instead of the using the original summaries. This was done to keep the experiments compatible with the LLMLingua Jiang et al. (2023); Pan et al. (2024) baselines.
@@
For the fine-tuning, a subset of the original SQuAD 2.0 Rajpurkar et al. (2016) dataset with $\sim$34k samples was used with 15 epochs of training. This subset was created by removing very long and very short contexts. We also focused on having same contexts with different questions to test out approach more rigorously. The test set contained $\sim$6k examples.
@@
The fine-tuning comprised training on a subset of the CodeSearchNet Husain et al. (2020) dataset with $\sim$25k samples for 4 epochs. Since code contexts need to be summarized differently that text ones, this subset was created by removing very long and very short code contexts from the original corpus. This was done to allow the model to learn the important sections of the complete context rather than just tokens. We tested on a test subset with $\sim$1300 examples of python code context.
@@ Appendix D Token-wise Relevance Reward
LLMLingua-2 - CodeSearchNet 4.74 27.06 8.08 18.63 80.71
LLMLingua-2 - Py150 4.15 (-0.59) 27.10 (+0.04) 7.70 (-0.38) 18.38 (-0.25) 80.75 (+0.04)
LLMLingua 7.01 (+2.27) 31.35 (+4.29) 13.09 (+5.01) 23.35 (+4.72) 81.48 (+0.78)
@@
TACO-RL - with Rouge1 9.57 (+4.83) 34.97 (+7.91) 15.85 (+7.77) 26.20 (+7.57) 81.85 (+1.14)
TACO-RL - with Bleu 11.23 (+6.49) 35.66 (+8.60) 17.98 (+9.90) 27.36 (+8.73) 82.03 (+1.32)
@@
LLMLingua-2 - CodeSearchNet 2.84 23.72 5.54 16.08 80.25
LLMLingua-2 - Py150 2.82 (-0.02) 23.68 (-0.04) 5.24 (-0.30) 15.74 (-0.34) 80.27 (+0.02)
LLMLingua 3.81 (+0.97) 26.44 (+2.72) 8.27 (+2.72) 18.79 (+2.71) 80.72 (+0.46)
@@
TACO-RL - with Rouge1 8.13 (+5.29) 34.18 (+10.46) 14.55 (+9.01) 25.54 (+9.46) 81.80 (+1.55)
TACO-RL - with Bleu 9.87 (+7.03) 35.94 (+12.22) 17.97 (+12.43) 27.60 (+11.52) 82.18 (+1.93)
@@
LLMLingua-2 - CodeSearchNet 2.26 21.99 4.54 14.78 80.05
LLMLingua-2 - Py150 2.36 (+0.10) 22.37 (+0.38) 4.40 (-0.14) 14.86 (+0.08) 80.11 (+0.06)
LLMLingua 2.28 (+0.03) 23.64 (+1.65) 5.78 (+1.23) 16.41 (+1.64) 80.24 (+0.18)
@@
TACO-RL - with Rouge1 6.27 (+4.01) 33.19 (+11.20) 13.37 (+8.83) 24.58 (+9.80) 81.75 (+1.70)
TACO-RL - with Bleu 8.18 (+5.92) 35.49 (+13.50) 17.40 (+12.86) 27.43 (+12.65) 82.24 (+2.19)
@@
LLMLingua-2 - CodeSearchNet 1.97 21.01 3.99 14.21 79.96
LLMLingua-2 - Py150 1.88 (-0.09) 21.07 (+0.06) 3.83 (-0.16) 14.14 (-0.07) 79.98 (+0.02)
LLMLingua 1.45 (-0.52) 21.39 (+0.37) 4.27 (+0.29) 14.74 (+0.53) 79.94 (-0.02)
@@
TACO-RL - with Rouge1 4.98 (+3.01) 31.84 (+10.83) 12.03 (+8.04) 23.68 (+9.47) 81.63 (+1.67)
TACO-RL - with Bleu 7.02 (+5.05) 34.96 (+13.95) 16.90 (+12.91) 26.90 (+12.69) 82.16 (+2.20)
@@
LLMLingua-2 - CodeSearchNet 1.64 20.14 3.48 13.56 79.87
LLMLingua-2 - Py150 1.78 (+0.14) 20.69 (+0.55) 3.47 (-0.01) 13.95 (+0.39) 79.92 (+0.05)
LLMLingua 1.04 (-0.61) 19.93 (-0.21) 3.42 (-0.06) 13.70 (+0.14) 79.72 (-0.15)
@@
TACO-RL - with Rouge1 4.22 (+2.58) 30.96 (+10.82) 11.58 (+8.10) 23.00 (+9.44) 81.58 (+1.71)
TACO-RL - with Bleu 6.01 (+4.37) 34.31 (+14.17) 16.10 (+12.62) 26.36 (+12.80) 82.12 (+2.25)
Results with Original Prompts 9.80 33.83 15.79 25.39 81.69

### v3

@@ Appendix A Details of Experiments
For this dataset, we used target (reference) summaries generated using GPT-4 instead of the using the original summaries. This was done to keep the experiments compatible with the LLMLingua Jiang et al. (2023); Pan et al. (2024) baselines.
@@
For the fine-tuning, a subset of the original SQuAD 2.0 Rajpurkar et al. (2016) dataset with $\sim$34k samples was used with 15 epochs of training. This subset was created by removing very long and very short contexts. We also focused on having same contexts with different questions to test out approach more rigorously. The test set contained $\sim$6k examples. During evaluation, original answers present in the dataset were treated as reference answers.
@@
We fine-tuned our model on a curated subset of $\sim$25k samples from the CodeSearchNet Husain et al. (2020) dataset over 4 epochs. Recognizing the unique challenges in code context summarization, we preprocessed the dataset by eliminating extremely long and short code contexts. This strategic filtering enables the model to focus on learning the most informative sections of code contexts, rather than getting distracted by peripheral tokens or overly verbose implementations.
The evaluation was conducted on a test subset comprising $\sim$1300 Python code examples. Due to the lack of usable summaries in the dataset, reference summaries were generated using complete original prompts with GPT 3.5.
@@ Appendix D Token-wise Relevance Reward
Appendix E Prompt Compression Comparison For Code Summarization
E.1 Original Code Context
1def yixia_download(url, output_dir = ’.’, merge = True, info_only = False, **kwargs):
2 """wrapper"""
3 hostname = urlparse(url).hostname
4 if ’n.miaopai.com’ == hostname:
5 smid = match1(url, r’n\.miaopai\.com/media/([^.]+)’)
6 miaopai_download_by_smid(smid, output_dir, merge, info_only)
7 return
8 elif ’miaopai.com’ in hostname: #Miaopai
9 yixia_download_by_scid = yixia_miaopai_download
10 site_info = "Yixia Miaopai"
11 scid = match1(url, r’miaopai\.com/show/channel/([^.]+)\.htm’) or \
12 match1(url, r’miaopai\.com/show/([^.]+)\.htm’) or \
13 match1(url, r’m\.miaopai\.com/show/channel/([^.]+)\.htm’) or \
14 match1(url, r’m\.miaopai\.com/show/channel/([^.]+)’)
15 elif ’xiaokaxiu.com’ in hostname: #Xiaokaxiu
16 yixia_download_by_scid = yixia_xiaokaxiu_download
17 site_info = "Yixia Xiaokaxiu"
18 if re.match(r’http://v.xiaokaxiu.com/v/.+\.html’, url): #PC
19 scid = match1(url, r’http://v.xiaokaxiu.com/v/(.+)\.html’)
20 elif re.match(r’http://m.xiaokaxiu.com/m/.+\.html’, url): #Mobile
21 scid = match1(url, r’http://m.xiaokaxiu.com/m/(.+)\.html’)
22 else:
23 pass
Listing 1: Original Code Context
E.2 Compressed Context - LLMLingua-2
1 ’n.miaopai
2(url[^’
3 ’miaopai.com’
4 Miaopai
5 match1(url.com/show/channel/([^.]+)\.htm
6 match1[^
7 match1([^.]+)’)
8 ’xiaokaxiu.com’
9_download
10 "Yixia Xiaokaxiu"
11 re.match(r’http://v.xiaokaxiu.com/v/.+\.html’,
12 match1(url r’http://v.xiaokaxiu/v/(.+).html’)
13 re.match(r’http://m.xiaokaxiu.com/m/.+\.html’
14 match1(url, r’http://m.xiaokaxiu.com/m(
15 else pass’
Listing 2: Compressed Code by LLMLingua-2
E.3 Compressed Context - TACO-RL
1def yixia_download(url, output_dir = ’.’, merge = True, info_only = False, **kwargs):
2 """wrapper"""
3 hostname =
4 if ’n.miaopai.com’ == hostname:
5 smid =
6 return
7 elif ’miaopai.com’ in hostname: #Miaopai
8 "Yixia Miaopai"
9 scid
10 elif ’xiaokaxiu.com’
11 Xiaokaxiu
12 if re #PC
13 scid
14 elif
15+
16 pass
Listing 3: Compressed Code by TACO-RL
@@ 11 Xiaokaxiu
LLMLingua-2 - CodeSearchNet 18.30 51.57 22.50 37.94 90.27
LLMLingua-2 - Py150 19.90 (+1.60) 52.93 (+1.36) 24.35 (+1.85) 39.65 (+1.71) 90.61 (+0.34)
LLMLingua 23.58(+5.28) 55.23 (+3.66) 28.64 (+6.14) 43.41 (+5.47) 91.12 (+0.85)
@@
TACO-RL - with Rouge1 29.97 (+11.67) 59.32 (+7.75) 33.99 (+11.49) 47.45 (+9.51) 91.76 (+1.49)
TACO-RL - with Bleu 35.04 (+16.74) 61.26 (+9.69) 38.83 (+16.34) 50.77 (+12.84) 92.26 (+1.98)
@@
LLMLingua-2 - CodeSearchNet 11.98 44.95 15.83 32.07 89.06
LLMLingua-2 - Py150 12.11 (+0.14) 44.94 (-0.01) 15.81 (-0.02) 32.10 (+0.02) 89.10 (+0.04)
LLMLingua 13.40 (+1.42) 46.29 (+1.34) 18.21 (+2.40) 34.49 (+2.42) 89.33 (+0.28)
@@
TACO-RL - with Rouge1 22.77 (+10.79) 54.39 (+9.44) 27.34 (+11.50) 41.81 (+9.74) 90.83 (+1.77)
TACO-RL - with Bleu 28.81 (+16.84) 57.32 (+12.37) 33.99 (+18.16) 46.90 (+14.82) 91.56 (+2.50)
@@
LLMLingua-2 - CodeSearchNet 8.93 41.07 12.30 28.87 88.31
LLMLingua-2 - Py150 9.35 (+0.42) 41.62 (+0.55) 12.81 (+0.51) 29.21 (+0.34) 88.45 (+0.14)
LLMLingua 9.06 (+0.13) 41.51 (+0.44) 13.69 (+1.39) 30.15 (+1.28) 88.34 (+0.02)
@@
TACO-RL - with Rouge1 17.59 (+8.66) 50.87 (+9.80) 23.03 (+10.73) 38.48 (+9.61) 90.22 (+1.91)
TACO-RL - with Bleu 24.23 (+15.30) 54.89 (+13.82) 31.26 (+18.96) 44.61 (+15.74) 91.14 (+2.83)
@@
LLMLingua-2 - CodeSearchNet 7.31 38.88 10.56 27.21 87.87
LLMLingua-2 - Py150 7.76 (+0.46) 39.09 (+0.21) 11.00 (+0.45) 27.46 (+0.26) 88.00 (+0.13)
LLMLingua 6.57 (-0.73) 38.63 (-0.25) 11.13 (+0.58) 27.81 (+0.60) 87.74 (-0.13)
@@
TACO-RL - with Rouge1 14.15 (+6.84) 47.94 (+9.06) 20.20 (+9.65) 36.10 (+8.89) 89.76 (+1.89)
TACO-RL - with Bleu 21.13 (+13.82) 52.67 (+13.79) 29.14 (+18.58) 42.84 (+15.63) 90.81 (+2.94)
@@
LLMLingua-2 - CodeSearchNet 6.54 37.28 9.58 26.20 87.64
LLMLingua-2 - Py150 6.79 (+0.25) 37.63 (+0.35) 9.95 (+0.37) 26.49 (+0.29) 87.69 (+0.05)
LLMLingua 5.31 (-1.23) 36.62 (-0.66) 9.76 (+0.18) 26.38 (+0.18) 87.36 (-0.29)
@@
TACO-RL - with Rouge1 12.38 (+5.84) 46.07 (+8.79) 18.73 (+9.15) 34.93 (+8.73) 89.53 (+1.89)
TACO-RL - with Bleu 18.46 (+11.92) 50.99 (+13.71) 27.33 (+17.75) 41.34 (+15.14) 90.52 (+2.88)
Results with Original Prompts 87.89 92.87 88.69 91.25 98.61
@@
Metric Mean Std 95% CI P-value
BLEU 21.2461 0.0991 [21.1230, 21.3692] 1.0000
ROUGE1 55.2168 0.0728 [55.1265, 55.3071] 1.0000
ROUGE2 31.7949 0.0387 [31.7469, 31.8430] 1.0000
ROUGEL 41.9159 0.0309 [41.8775, 41.9542] 1.0000
BERTScore 90.9349 0.0060 [90.9274, 90.9424] 1.0000
Table 8: One-sample t-test results for 2x compression over 5 runs on the MeetingBank Dataset for TACO-RL. The results show mean scores, standard deviation, 95% confidence intervals, and p-values for various evaluation metrics.
Method BLEU ROUGE1 ROUGE2 ROUGEL BERTScore
LLMlingua-2 - Wikitext 16.71 52.58 27.73 39.05 90.47
Original Prompts 21.50 55.19 33.03 42.90 91.12
TACO-RL - with entropy term 21.35 55.34 31.88 42.17 90.95
TACO-RL - without entropy term 17.62 53.39 28.59 39.49 90.57
Table 9: Performance comparison of TACO-RL with and without the entropy term $\lambda H(\mathbf{p})$ on MeetingBank dataset at 2x compression ratio. Best results are shown in bold.